OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (agent47):

y=x+9

OpenStudy (agent47):

that would be point slope form.

OpenStudy (anonymous):

I'm sorry, that's the wrong equation. It's y-3=1/2(x+6)

OpenStudy (agent47):

y=(1/2)x+6

OpenStudy (agent47):

That's point slope form.

OpenStudy (anonymous):

Ok, thanks

OpenStudy (precal):

yes but I believe you are also, looking for standard form

OpenStudy (precal):